### v1

# Ando type dilation for completely contractive covariant representations

Abstract.
Solel proved that every completely contractive covariant representation of a product system $\mathbb{E}$ over $\mathbb{N}^{2}_{0}$ can be dilated to an isometric representation of $\mathbb{N}^{2}_{0},$ which covers Ando’s dilation result when $C^{*}$-algebra $\mathcal{M}=\mathbb{C}$ and $\mathbb{E}(\mathbf{n})=\mathbb{C},$ where $\mathbf{n}\in\mathbb{N}^{2}_{0}.$ That arises a natural question that exactly which type of representations dilate completely contractive covariant representation of product system $\mathbb{E}$ over $\mathbb{N}^{2}_{0}?$ To identify the isometric covariant representations for which the above has an affirmative answer is the main object of this article.
Key words and phrases:
Hilbert $C^{*}$-modules, Covariant representations, Invariant subspaces, Tensor product
1991 Mathematics Subject Classification:
Primary 46L08; Secondary 47A13, 47A15, 47L55
1. Introduction
The theory of dilation is a fundamental result established by Sz.-Nagy [14], which asserts that any contraction on a Hilbert space can be dilated to an isometry acting on a larger Hilbert space. This result laid the foundation for a deeper understanding of operator theory, providing a powerful tool for studying contractions through their isometric extensions. Building on Sz.-Nagy’s foundational work, Ando [1] significantly advanced the theory by constructing isometric dilations for pairs of commuting contractions. That is, for any pair of commuting contractions $(T_{1},T_{2})$ on a Hilbert space $\mathcal{H},$ then there exists a pair of commuting isometries $(V_{1},V_{2})$ on a Hilbert space $\mathcal{K}\supseteq\mathcal{H}$ such that
$T_{1}^{n_{1}}T_{2}^{n_{2}}=P_{\mathcal{H}}V_{1}^{n_{1}}V_{2}^{n_{2}}|_{\mathcal{H}}\quad\quad\mbox{for all}\quad\quad(n_{1},n_{2})\in\mathbb{Z}_{+}^{2},$
where $P_{\mathcal{H}}$ denotes the orthogonal projection onto ${\mathcal{H}}.$ Ando’s theorem thus confirms that every pair of commuting contractions can be dilated to a pair of commuting isometries, providing a complete solution in the two-variable case. Parrott [15] gave a counterexample showing that for $n\geq 3,$ $n$-tuples of commuting contractions do not generally admit isometric dilations. This striking limitation highlights the intrinsic difficulties of multivariable dilation theory and motivates a more nuanced investigation into the specific conditions under which such dilations might still be possible.
In recent years, a growing body of literature has explored the interplay between dilation and positivity, revealing that certain classes of operator tuples under specific positivity assumptions can indeed possess isometric dilations. See [2, 4, 5, 8, 13] for more details in the polydisc setup. The main motivation behind our interest comes from a recent work by Das and Sarkar [5], they obtained an explicit isometric dilation of pair of commuting contractions in the following theorem:
Theorem 1.1.
Let $(T_{1},T_{2})$ be a pair of commuting contractions on a Hilbert space $\mathcal{H}.$ Let $T_{1}$ be pure and dim $\mathcal{D}_{T_{j}}<\infty,j=1,2.$ Then there exist an isometry $\Pi:\mathcal{H}\to H^{2}_{\mathcal{D}_{T_{1}}}(\mathbb{D})$ and an inner function $\psi\in H^{\infty}_{\mathcal{B}(\mathcal{D}_{T_{1}})}(\mathbb{D})$ such that
$\Pi T_{1}^{*}=M_{z}^{*}\Pi\quad\quad\mbox{and}\quad\quad\Pi T_{2}^{*}=M_{\psi}^{*}\Pi,$
where $\mathcal{D}_{T_{j}}=\overline{ran}(I_{\mathcal{H}}-T_{j}T^{*}_{j})$ for $j=1,2.$
Operator and function theories are closely intertwined and have long been regarded as well-established fields with significant connections to a wide range of disciplines. These areas continue to advance, both by addressing traditional problems and by extending classical ideas to more generalized frameworks. One such extension is the concept of Hilbert $C^{*}$-modules, developed through $C^{*}$-correspondences and covariant representations.
Cuntz [3] studied a $C^{*}$-algebra, known as Cuntz algebra, generated by isometries $V_{1},\dots,V_{n}$ $(n\geq 2)$ on a Hilbert space with $V_{i}$’s have orthogonal range. Wold decomposition for two isometries with orthogonal range was first introduced by Frazho [7]. Popescu [18] extended this decomposition to the case of an infinite sequence of isometries with orthogonal final spaces, and explored a minimal isometric dilation for an infinite sequence of noncommuting contractive operators on a Hilbert space. Pimsner [17] generalized the construction of Cuntz algebras with the help of the notion of isometric representations of $C^{*}$-correspondences. Muhly and Solel [11] discussed the Wold decomposition for the isometric covariant representations of $C^{*}$-correspondences based on [18].
Muhly and Solel [12] presented isometric dilation for completely contractive covariant representations of $C^{*}$-correspondences. Solel [24] proved that every completely contractive representations of product system of $C^{*}$-correspondences over the semigroup $\mathbb{N}_{0}^{2}$ can be dilated to isometric representations. Solel [25] studied completely contractive representations of product system of $C^{*}$-correspondences over $\mathbb{N}_{0}^{k},$ and gave a necessary and sufficient condition for such a representation to have a regular isometric dilation. In [22] Skalski established some sufficient conditions for the existence of a not necessarily regular isometric dilation of a completely contractive representation of the product system and discussed the difference between regular and *-regular dilations. He also analysed the case of product systems associated with higher-rank graphs. The main purpose of this paper is to follow this direction along [5] and to investigate which type of representation dilates completely contractive covariant representations of product system.
1.1. Notations and Preliminaries
We begin this section by recalling few concepts and their properties about $C^{*}$-correspondences from [10, 12, 16, 19, 21]. Let $E$ be a Hilbert $C^{*}$-module over a $C^{*}$-algebra $\mathcal{M}.$ By $B^{a}(E)$ we denote the $C^{*}$-algebra of all adjointable operators on $E$. We say that the module $E$ is a $C^{*}$-correspondence over $\mathcal{M}$ if there exists a left $\mathcal{M}$-module structure through a non-zero *-homomorphism $\phi:\mathcal{M}\to B^{a}(E)$ in the following sense
$a\xi:=\phi(a)\xi\quad\quad(a\in\mathcal{M},\xi\in E).$
In this article, each *-homomorphism considered is essential, this means that, the closed linear span of $\phi(\mathcal{M})E$ equals $E.$ If $F$ is another $C^{*}$-correspondence over $\mathcal{M},$ then we may consider the notion of tensor product $F\bigotimes_{\phi}E$ (cf. [10]) which satisfy
$(\eta_{1}a)\otimes\xi_{1}=\eta_{1}\otimes\phi(a)\xi_{1},$
$\langle\eta_{1}\otimes\xi_{1},\eta_{2}\otimes\xi_{2}\rangle=\langle\xi_{1},\phi(\langle\eta_{1},\eta_{2}\rangle)\xi_{2}\rangle$
for every $\eta_{1},\eta_{2}\in F,$ $\xi_{1},\xi_{2}\in E$ and $a\in\mathcal{M}.$ In this paper, we adopt the following notations: $\mathcal{H}$ denotes a Hilbert space, $E$ represents a $C^{*}$-correspondence over $\mathcal{M}$, and $B(\mathcal{H})$ refers to the algebra of all bounded linear operators on $\mathcal{H}.$ The notion of completely contractive covariant representations plays a crucial role in this paper.
Definition 1.2.
Let $\sigma:\mathcal{M}\to B(\mathcal{H})$ be a representation and $T:E\to B(\mathcal{H})$ be a linear map. Then the tuple $(\sigma,T)$ is said to be a covariant representation of $E$ on $\mathcal{H}$ if
$T(a\xi a^{\prime})=\sigma(a)T(\xi)\sigma(a^{\prime})\quad\quad(\xi\in E,a,a^{\prime}\in\mathcal{M}).$
We say that the covariant representation $(\sigma,T)$ is completely bounded (respectively, completely contractive) if $T$ is completely bounded (respectively, completely contractive ) as a linear map on $E$ (endowed with the usual operator space structure by viewing as a left corner in the linking algebra $\mathcal{L}_{E}$ of $E$). Moreover, $(\sigma,T)$ is called isometric if
$T(\xi)^{*}T(\zeta)=\sigma(\langle\xi,\zeta\rangle)\quad\quad(\xi,\zeta\in E).$
Muhly and Solel gave the following key lemma in [12] which is useful to classify the completely bounded covariant representations of a $C^{*}$-correspondence:
Lemma 1.3.
The map $(\sigma,T)\mapsto\widetilde{T}$ provides a bijection between the collection of all completely bounded covariant (respectively, completely contractive covariant) representations $(\sigma,T)$ of $E$ on $\mathcal{H}$ and the collection of all bounded (respectively, contractive) linear maps $\widetilde{T}:~{}\mbox{$E\bigotimes_{\sigma}\mathcal{H}\to\mathcal{H}$}$ defined by
$\widetilde{T}(\xi\otimes h):=T(\xi)h\quad\quad(\xi\in E,h\in\mathcal{H}),$
such that $\widetilde{T}(\phi(a)\otimes I_{\mathcal{H}})=\sigma(a)\widetilde{T}$, $a\in\mathcal{M}$. Moreover, $\widetilde{T}$ is isometry (respectively, co-isometry) if and only if $(\sigma,T)$ is isometric (respectively, co-isometric).
Definition 1.4.
Suppose that $(\sigma,T)$ is a completely bounded covariant representation of ${E}$ on $\mathcal{H}.$ A closed subspace $\mathcal{W}$ of $\mathcal{H}$ is said to be $(\sigma,T)$-invariant $(resp.(\sigma,T)$-reducing) (cf. [21]) if it is $\sigma(\mathcal{M})$-invariant and (resp. both $\mathcal{W},\mathcal{W}^{\perp}$) is invariant by each operator $T(\xi)$ for all $\xi\in E.$ The restriction provides a new representation $(\sigma,T)|_{\mathcal{W}}$ of $E$ on $\mathcal{W}.$
A completely bounded covariant representation $(\sigma,T)$ of $E$ on $\mathcal{H}$ is said to be completely non unitary if there doesn’t exists a non-zero reducing subspace $\mathcal{W}$ of $\mathcal{H}$ such that $(\sigma,T)|_{\mathcal{W}}$ is isometric as well as co-isometric representation.
For each $n\in\mathbb{N},$ $E^{\otimes n}:=E\otimes_{\phi}\cdots\otimes_{\phi}E$ ($n$ times) ($E^{\otimes 0}=\mathcal{M}$) is a $C^{*}$-correspondence over the $C^{*}$-algebra $\mathcal{M}$, where the left action of $\mathcal{M}$ on $E^{\otimes n}$ is given by
$\phi^{n}(a)(\xi_{1}\otimes\cdots\otimes\xi_{n}):=\phi(a)\xi_{1}\otimes\cdots\otimes\xi_{n}.$
For $n\in\mathbb{N},$ define $\widetilde{T}_{n}:E^{\otimes n}\otimes\mathcal{H}\to\mathcal{H}$ by
$\widetilde{T}_{n}(\xi_{1}\otimes\dots\otimes\xi_{n}\otimes h)=T(\xi_{1})\dots T(\xi_{n})h,\quad\xi_{i}\in E,h\in\mathcal{H}.$
The Fock module $\mathcal{F}(E)=\bigoplus_{n\geq 0}E^{\otimes n}$ is a $C^{*}$-correspondence over a $C^{*}$-algebra $\mathcal{M}$, with left action of $\mathcal{M}$ on $\mathcal{F}(E)$ is given by
$\phi_{\infty}(a)(\oplus_{n\geq 0}\xi_{n})=\oplus_{n\geq 0}a\cdot\xi_{n},\>\>\xi_{n}\in E^{\otimes n},a\in\mathcal{M}.$
Let $\xi\in E$, we define the creation operator $T_{\xi}$ on $\mathcal{F}(E)$ by
$T_{\xi}(\eta):=\xi\otimes\eta,\>\>\eta\in E^{\otimes n}.$
Let $\pi$ be a representation of $\mathcal{M}$ on $\mathcal{H}$. Define an isometric covariant representation $(\rho,S)$ of $E$ on a Hilbert space $\mathcal{F}(E)\otimes_{\pi}\mathcal{H}$ by
$\rho(a):$ $=\phi_{\infty}(a)\otimes I_{\mathcal{H}}\>\>,a\in\mathcal{M}$
$S(\xi):$ $=T_{\xi}\otimes I_{\mathcal{H}},\>\>\xi\in E,$
such a representation $(\rho,S)$ is referred to as the induced representation (cf. [23]) induced by $\pi.$ Let $(\sigma,T)$ and $(\psi,V)$ be two completely bounded covariant representations of $E$ on the Hilbert spaces $\mathcal{H}$ and $\mathcal{K},$ respectively. We say that $(\sigma,T)$ is isomorphic to $(\psi,V)$ if there exists a unitary operator $U:\mathcal{H}\to\mathcal{K}$ such that $U\sigma(a)=\psi(a)U$ and $UT(\xi)=V(\xi)U$ for all $\xi\in E,a\in\mathcal{M}.$ A closed subspace $\mathcal{W}\subseteq\mathcal{H}$ is called wandering subspace for $(\sigma,T)$ if $\mathcal{W}$ is orthogonal to $\widetilde{T}_{n}(E^{\otimes n}\otimes\mathcal{W})$ for all $n\in\mathbb{N}.$ We say that the wandering subspace $\mathcal{W}$ is generating for $(\sigma,T)$ if
$\mathcal{H}=\bigvee_{n\geq 0}\widetilde{T}_{n}(E^{\otimes n}\otimes\mathcal{W}).$
The following theorem is an analogue of [14, Theorem 3.2].
Theorem 1.5.
Let $(\sigma,T)$ be a completely contractive covariant representation of $E$ on $\mathcal{H}.$ Then there exists a unique orthogonal decomposition of $\mathcal{H}=\mathcal{H}_{0}\oplus\mathcal{H}_{1}$ such that the following holds:
  1. (1)
    $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ reduce $(\sigma,T).$
  2. (2)
    $(\sigma,T)|_{\mathcal{H}_{0}}$ is an isometric as well as co-isometric representation.
  3. (3)
    $(\sigma,T)|_{\mathcal{H}_{1}}$ is completely non unitary.
Proof.
We define
$\mathcal{H}_{0}:$ =$=\{h\in\mathcal{H}:\|(I_{E^{\otimes n}}\otimes\widetilde{T}_{m})(\xi_{n+m}\otimes h)\|=\|\xi_{n+m}\otimes h\|\quad\mbox{and}\quad$
$\|(I_{E^{\otimes n}}\otimes\widetilde{T}_{m}^{*})(\xi_{n}\otimes h)\|=\|\xi_{n}\otimes h\|\quad\mbox{for}\quad n,m\in\mathbb{N}_{0},\xi_{n+m}\in E^{\otimes n+m}\}.$
It is easy to verify that $\mathcal{H}_{0}$ is a closed subspace of $\mathcal{H}.$ Now we need to show that $\mathcal{H}_{0}$ reduces $(\sigma,T).$ Let $h\in\mathcal{H}_{0},$ we want to prove that $\widetilde{T}(\xi\otimes h)\in\mathcal{H}_{0}$ for all $\xi\in E.$ Note that
$\|(I_{E^{\otimes n}}\otimes\widetilde{T}_{m})(\xi_{n+m}\otimes\widetilde{T}(\xi\otimes h))\|=\|(I_{E^{\otimes n}}\otimes\widetilde{T}_{m+1})(\xi_{n+m}\otimes\xi\otimes h)\|$
$=\|\xi_{n+m}\otimes\xi\otimes h\|=\|(I_{E^{\otimes n+m}}\otimes\widetilde{T})(\xi_{n+m}\otimes\xi\otimes h)\|=\|\xi_{n+m}\otimes\widetilde{T}(\xi\otimes h)\|.$
If $\|\widetilde{T}(\xi\otimes h)\|=\|\xi\otimes h\|,$ then $\widetilde{T}^{*}\widetilde{T}(\xi\otimes h)=\xi\otimes h$ we have
$\|(I_{E^{\otimes n}}\otimes\widetilde{T}_{m}^{*})(\xi_{n}\otimes\widetilde{T}(\xi\otimes h))\|$
$=\|(I_{E^{\otimes n}}\otimes(I_{E}\otimes\widetilde{T}_{m-1}^{*})\widetilde{T}^{*})(\xi_{n}\otimes\widetilde{T}(\xi\otimes h))\|$
$=\|\xi_{n}\otimes(I_{E}\otimes\widetilde{T}_{m-1}^{*})\widetilde{T}^{*}\widetilde{T}(\xi\otimes h)\|=\|\xi_{n}\otimes(I_{E}\otimes\widetilde{T}_{m-1}^{*})(\xi\otimes h)\|$
$=\|(I_{E^{\otimes n+1}}\otimes\widetilde{T}_{m-1}^{*})(\xi_{n}\otimes\xi\otimes h)\|=\|\xi_{n}\otimes\xi\otimes h\|$
$=\|(I_{E^{\otimes n+1}}\otimes\widetilde{T})(\xi_{n}\otimes\xi\otimes h)\|=\|\xi_{n}\otimes\widetilde{T}(\xi\otimes h)\|.$
This implies that $\widetilde{T}(\xi\otimes h)\in\mathcal{H}_{0}.$ Also we observe that
$\|(I_{E^{\otimes n}}\otimes\widetilde{T}_{m})(\xi_{n+m-1}\otimes\widetilde{T}^{*}h)\|=\|\xi_{n+m-1}\otimes\widetilde{T}^{*}h\|,$
and
$\|(I_{E^{\otimes n}}\otimes\widetilde{T}_{m}^{*})(\xi_{n-1}\otimes\widetilde{T}^{*}h)\|=\|\xi_{n-1}\otimes\widetilde{T}^{*}h\|,\quad\mbox{for all}\quad m,n\geq 1.$
This implies that $\mathcal{H}_{0}$ reduces $(\sigma,T).$ Also it is easy to see that $(\sigma,T)|_{\mathcal{H}_{0}}$ is an isometric as well as co-isometric representation. Since $\mathcal{H}_{1}=\mathcal{H}\ominus\mathcal{H}_{0},$ $(\sigma,T)|_{\mathcal{H}_{1}}$ is completely non-unitary. Indeed, let $\mathcal{K}$ be a subspace of $\mathcal{H}_{1}$ such that $(\sigma,T)|_{\mathcal{K}}$ is a unitary. It follows that $\mathcal{K}\subseteq{\mathcal{H}_{0}},$ which is a contradiction.
Uniqueness: Let $\mathcal{H}=\mathcal{H}^{\prime}_{0}\oplus\mathcal{H}^{\prime}_{1}$ be an orthogonal decomposition of $\mathcal{H}$ such that $(\sigma,T)|_{\mathcal{H}^{\prime}_{0}}$ is an isometric as well as co-isometric representation, then $\mathcal{H}^{\prime}_{0}\subseteq\mathcal{H}_{0}.$ The spaces $\mathcal{H}^{\prime}_{0}$ and $\mathcal{H}_{0}$ reduce $(\sigma,T),$ and hence $(\sigma,T)|_{\mathcal{H}_{0}\ominus\mathcal{H}^{\prime}_{0}}$ is an isometric as well as co-isometric representation. Note that $\mathcal{H}_{0}\ominus\mathcal{H}^{\prime}_{0}\subseteq\mathcal{H}\ominus\mathcal{H}^{\prime}_{0}=\mathcal{H}^{\prime}_{1}.$ Since $(\sigma,T)|_{\mathcal{H}^{\prime}_{1}}$ is completely non unitary, $\mathcal{H}_{0}\ominus\mathcal{H}^{\prime}_{0}=\{0\}.$ Therefore $\mathcal{H}_{0}=\mathcal{H}^{\prime}_{0},$ and hence the orthogonal decomposition is unique. ∎
Let $(\sigma,T)$ be a completely contractive covariant representation of $E$ on $\mathcal{H},$ that is, $I_{\mathcal{H}}-\widetilde{T}\widetilde{T}^{*}\geq 0.$ Define the positive operator $\Delta_{*}(T):=(I-\widetilde{T}{\widetilde{T}}^{*})^{\frac{1}{2}}$ and defect space $\mathcal{D}_{*,T}:=\overline{Im\Delta_{*}(T)}$. Since $\widetilde{T}\widetilde{T}^{*}\sigma(a)=\widetilde{T}(\phi(a)\otimes I_{\mathcal{H}})\widetilde{T}^{*}=\sigma(a)\widetilde{T}\widetilde{T}^{*}$ for all $a\in\mathcal{M},$ we get $\Delta_{*}(T)\in\sigma(\mathcal{M})^{\prime}.$ Thus $\mathcal{D}_{*,T}$ is $\sigma(\mathcal{M})$-reducing. Define $(\sigma_{0},T_{0}):=(\sigma,T)|_{\mathcal{D}_{*,T}},$ for simplicity we write $(\sigma,T)$ instead of $(\sigma_{0},T_{0})$. The Poisson kernel of $T$ is the operator $\Pi:\mathcal{H}\longrightarrow\mathcal{F}(E)\otimes_{\sigma}\mathcal{D}_{*,T}$ defined by
$\Pi h:=\sum_{n\geq 0}(I_{E^{\otimes n}}\otimes\Delta_{*}(T))\widetilde{T}_{n}^{*}h,\>\>\>h\in\mathcal{H}.$
Now we recall an isometric dilation of completely contractive covariant representations of a $C^{*}$-correspondence due to Muhly and Solel [12].
Theorem 1.6.
Let $(\sigma,T)$ be a completely contractive covariant representation of $E$ on $\mathcal{H}.$ Then the poisson kernel $\Pi$ is a contraction and
$\Pi\sigma(b)=\rho(b)\Pi,\>\>b\in\mathcal{M}\quad\mbox{and}\>\>(I_{E}\otimes\Pi)\widetilde{T}^{*}=\widetilde{S}^{*}\Pi.$
We say that the induced representation $(\rho,S)$ of $E$ induced by $\sigma|_{\mathcal{D}_{*,T}}$ is an isometric dilation of $(\sigma,T).$ Moreover, $\Pi$ is an isometry if and only if $(\sigma,T)$ is pure (that is, SOT-$\lim_{n\to\infty}\widetilde{T}_{n}\widetilde{T}_{n}^{*}=0$).
1.2. Completely contractive representations over product system
The central tool we require a product system of $C^{*}$-correspondences (see [6, 21, 23, 25]): The product system $\mathbb{E}$ is defined by a family of $C^{*}$-correspondences $\{E_{1},\ldots,E_{k}\},k\in\mathbb{N}$ and by the unitary isomorphisms $t_{i,j}:E_{i}\otimes E_{j}\to E_{j}\otimes E_{i}$ ($i>j$). Using these identifications, for all ${\bf n}=(n_{1},\cdots,n_{k})\in\mathbb{N}_{0}^{k}$ the correspondence $\mathbb{E}({\bf n})$ is identified with $E_{1}^{\otimes^{n_{1}}}\otimes\cdots\otimes E_{k}^{\otimes^{n_{k}}}.$ We use notations $t_{i,i}={\textrm{id}}_{E_{i}\otimes E_{i}}$ and $t_{i,j}=t_{j,i}^{-1}$ if $i<j.$
Definition 1.7.
Assume $\mathbb{E}$ to be a product system over $\mathbb{N}_{0}^{k}$.
  • (1)
    Let $\sigma$ be a representation of $\mathcal{M}$ on $\mathcal{H},$ and let $T^{(i)}:E_{i}\to B(\mathcal{H})$ be a linear maps for $1\leq i\leq k.$ The tuple $(\sigma,T^{(1)},\dots,T^{(k)})$ is called a completely bounded (respectively, completely contractive) covariant representation of a product system $\mathbb{E}$ on $\mathcal{H}$ if each pair $(\sigma,T^{(i)})$ is a completely bounded (respectively, completely contractive) covariant representation of $E_{i}$ on $\mathcal{H}$ and satisfy the commutation relation
    $\widetilde{T}^{(i)}(I_{E_{i}}\otimes\widetilde{T}^{(j)})=\widetilde{T}^{(j)}(I_{E_{j}}\otimes\widetilde{T}^{(i)})(t_{i,j}\otimes I_{\mathcal{H}})\quad\quad\mbox{for}\quad 1\leq i,j\leq k.$
  • (2)
    The covariant representation $(\sigma,T^{(1)},\ldots,T^{(k)})$ is called isometric (respectively, co-isometric) if each $(\sigma,T^{(i)})$ is an isometric (respectively, co-isometric) representation of $E_{i}.$ The representation $(\sigma,T^{(1)},\ldots,T^{(k)})$ is called pure if each $(\sigma,T^{(i)})$ is pure representation of $E_{i}$ on $\mathcal{H}.$
For each $1\leq i\leq k$ and $l\in\mathbb{N},$ define $\widetilde{T}^{(i)}_{l}:E_{i}^{\otimes l}\otimes\mathcal{H}\to\mathcal{H}$ by
$\widetilde{T}^{(i)}_{l}(\xi_{1}\otimes\cdots\otimes\xi_{l}\otimes h):=T^{(i)}(\xi_{1})\cdots T^{(i)}(\xi_{l})h$
for all $\xi_{1},\ldots,\xi_{l}\in E_{i},h\in\mathcal{H}.$ It is easy to verify that
$\widetilde{T}^{(i)}_{l}=\widetilde{T}^{(i)}(I_{E_{i}}\otimes\widetilde{T}^{(i)})\cdots(I_{E^{\otimes l-1}_{i}}\otimes\widetilde{T}^{(i)}).$
For $\mathbf{n}=(n_{1},\cdots,n_{k})\in\mathbb{N}_{0}^{k}$, we use notation $\widetilde{T}_{\mathbf{n}},$ where $\widetilde{T}_{\mathbf{n}}:\mathbb{E}(\mathbf{n})\otimes_{\sigma}\mathcal{H}\to\mathcal{H}$ is given by
$\widetilde{T}_{\mathbf{n}}=\widetilde{T}_{n_{1}}^{(1)}\left(I_{E_{1}^{\otimes n_{1}}}\otimes\widetilde{T}_{n_{2}}^{(2)}\right)\cdots\left(I_{E_{1}^{\otimes n_{1}}\otimes\cdots\otimes E_{k-1}^{\otimes{n_{k-1}}}}\otimes\widetilde{T}_{n_{k}}^{(k)}\right).$
Using Lemma 1.3, let us define the bi-module map $T_{\mathbf{n}}:\mathbb{E}(\mathbf{n})\to B(\mathcal{H})$ by
$T_{\mathbf{n}}(\xi)h=\widetilde{T}_{\mathbf{n}}(\xi\otimes h)\quad\mbox{for}\quad~{}\xi\in\mathbb{E}(\mathbf{n}),h\in\mathcal{H}.$
Define the Fock module of $\mathbb{E}$ over $\mathbb{N}^{k}_{0}$ by $\mathcal{F}(\mathbb{E}):=\bigoplus_{\mathbf{n}\in\mathbb{N}^{k}_{0}}\mathbb{E}(\mathbf{n}).$ Note that $\mathcal{F}(\mathbb{E})$ is a $C^{*}$-correspondence with the left action $\phi_{\infty}$ of $\mathcal{M}$ on $\mathcal{F}(\mathbb{E})$ given by
$\phi_{\infty}(b)(\bigoplus_{{\bf{n}}\in\mathbb{N}_{0}^{k}}\xi_{\bf{n}})=\bigoplus_{{\bf{n}}\in\mathbb{N}_{0}^{k}}b\xi_{\bf{n}},\>b\in\mathcal{M},\xi_{\bf{n}}\in\mathbb{E}(\bf{n}).$
Let $\pi$ be a representation of $\mathcal{M}$ on a Hilbert space $\mathcal{K}.$ For $i\in\{1,\cdots,k\},$ define an isometric covariant representation $(\rho,S^{(i)})$ of $E_{i}$ on the Hilbert space $\mathcal{F}(\mathbb{E})\otimes_{\pi}\mathcal{K}$ (cf. [23]) by
$\rho(a):=\phi_{\infty}(a)\otimes I_{\mathcal{K}},\>\>a\in\mathcal{M}\quad\mbox{and}\quad S^{(i)}(\xi_{i})=T_{{\xi}_{i}}\otimes I_{\mathcal{K}},\>\xi_{i}\in E_{i},$
where $T_{\xi_{i}}$ denotes the creation operator on $\mathcal{F}(\mathbb{E})$ determined by $\xi_{i}.$ We say that $(\rho,S^{(1)},\dots,S^{(k)})$ is an induced representation of $\mathbb{E}$ over $\mathbb{N}^{k}_{0}$ induced by $\pi.$
Suppose $u=\{u_{1},u_{2},\dots,u_{r}\}\subseteq\{1,2,\dots,k\}$ such that $u_{1}<u_{2}<\dots<u_{r}.$ Define $\mathbf{e}(u):=\mathbf{e}_{u_{1}}+\mathbf{e}_{u_{2}}+\dots+\mathbf{e}_{u_{r}},$ where $\mathbf{e}_{i}:=(0,0,\dots,1,0,\dots 0),$ i.e., $1$ at $i^{th}$ place and 00 otherwise. Next we define the following Brehmer-Solel condition:
Definition 1.8.
Let $(\sigma,T^{(1)},T^{(2)},\dots,T^{(k)})$ be a completely contractive representation of $\mathbb{E}$ over $\mathbb{N}^{k}_{0}$ on $\mathcal{H}.$ We say that it satisfies Brehmer-Solel condition (cf. [25]) if it satisfies the condition
$\sum_{u\subseteq\{1,2,\dots,k\}}(-1)^{|u|}\widetilde{V}_{\mathbf{e}(u)}\widetilde{V}_{\mathbf{e}(u)}^{*}\geq 0.$
The following theorem gave an isometric dilation of completely contractive covariant representations of product system due to Rohilla, Trivedi and Veerabathiran [19, Theorem 3.5]:
Theorem 1.9.
Let $(\sigma,T^{(1)},T^{(2)},\dots,T^{(k)})$ be a pure completely contractive covariant representation of $\mathbb{E}$ on $\mathcal{H},$ satisfying Brehmer-Solel condition. Then there exist a Hilbert space $\mathcal{D}_{*,k}(T)$ and an isometry $\Pi:\mathcal{H}\to\mathcal{F}(\mathbb{E})\otimes_{\sigma}\mathcal{D}_{*,k}(T)$ such that
$\Pi\sigma(a)=\rho(a)\Pi\>\>\>\>\mbox{and}\>\>\>\>(I_{E_{i}}\otimes\Pi)\widetilde{T}^{(i)*}=\widetilde{S}^{(i)*}\Pi\quad\mbox{for}\quad a\in\mathcal{M},1\leq i\leq k.$
We say that the induced representation $(\rho,S^{(1)},\dots,S^{(k)})$ of $\mathbb{E}$ induced by $\sigma|_{\mathcal{D}_{*,T}}$ is an isometric dilation of $(\sigma,T^{(1)},\dots,T^{(k)}).$
2. Ando type dilation for completely contractive representations of product system over $\mathbb{N}^{2}_{0}$
This section aims to construct an explicit isometric dilation of completely contractive covariant representations of product system over $\mathbb{N}_{0}^{2}.$ In particular, the isometric dilation of completely contractive covariant representation with finite defect indices, as this will play a key role in the subsequent analysis.
Assume $\mathbb{E}$ to be a product system over $\mathbb{N}^{2}_{0}$. Let $(\sigma,T^{(1)},T^{(2)})$ be a completely contractive covariant representation of $\mathbb{E}$ on $\mathcal{H}.$ Define $\widetilde{T}:=\widetilde{T}^{(1)}(I_{E_{1}}\otimes\widetilde{T}^{(2)}),$ then by Lemma 1.3 $(\sigma,T)$ is also a completely contractive covariant represenation of $E_{1}\otimes E_{2}$ on $\mathcal{H}.$ Note that
$I_{\mathcal{H}}-\widetilde{T}^{(1)}\widetilde{T}^{(1)^{*}}+\widetilde{T}^{(1)}(I_{E_{1}}\otimes(I_{\mathcal{H}}-\widetilde{T}^{(2)}\widetilde{T}^{(2)^{*}}))\widetilde{T}^{(1)^{*}}$
$=I_{\mathcal{H}}-\widetilde{T}^{(1)}(I_{E_{1}}\otimes\widetilde{T}^{(2)})(I_{E_{1}}\otimes\widetilde{T}^{(2)^{*}})\widetilde{T}^{(1)^{*}}=I_{\mathcal{H}}-{\widetilde{T}}{\widetilde{T}^{*}}$
$=I_{\mathcal{H}}-\widetilde{T}^{(2)}\widetilde{T}^{(2)^{*}}+\widetilde{T}^{(2)}(I_{E_{2}}\otimes(I_{\mathcal{H}}-\widetilde{T}^{(1)}\widetilde{T}^{(1)^{*}}))\widetilde{T}^{(2)^{*}}.$
This implies that, for $h\in\mathcal{H}$
$\|\Delta_{*}(T^{(1)})h\|^{2}+\|(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}h\|^{2}$
$=\|(I_{E_{2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}}h\|^{2}+\|\Delta_{*}(T^{(2)})h\|^{2}.$
Therefore we get an isometry
$U:\{\Delta_{*}(T^{(1)})h\oplus(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}h:~{}h\in\mathcal{H}\}\to$
$\{(I_{E_{2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}}h\oplus\Delta_{*}(T^{(2)})h:~{}h\in\mathcal{H}\}$
defined by
$U(\Delta_{*}(T^{(1)})h,(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}h)$
(2.1) $=((I_{E_{2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}}h,\Delta_{*}(T^{(2)})h),$
$h\in\mathcal{H}.$ Let $\phi_{1}$ and $\phi_{2}$ be the left actions of $\mathcal{M}$ on $E_{1}$ and $E_{2},$ respectively. It is easy to check that for $a\in\mathcal{M}$ and $h\in\mathcal{H},$ $U$ is a module map, that is,
$U\rho^{{}^{\prime}}(a)=\rho^{{}^{\prime\prime}}(a)U,$
where $\rho^{{}^{\prime}}(a)=\big{(}\sigma(a)|_{\mathcal{D}_{*,T^{(1)}}},\phi_{1}(a)\otimes I_{\mathcal{D}_{*,T^{(2)}}}\big{)}$ and $\rho^{{}^{\prime\prime}}(a)=\big{(}\phi_{2}(a)\otimes I_{\mathcal{D}_{*,T^{(1)}}},\\ \sigma(a)|_{\mathcal{D}_{*,T^{(2)}}}\big{)}$ are the left actions of $\mathcal{M}$ on the Hilbert spaces $\mathcal{D}_{*,T^{(1)}}\oplus(E_{1}\otimes\mathcal{D}_{*,T^{(2)}})$ and $(E_{2}\otimes\mathcal{D}_{*,T^{(1)}})\oplus\mathcal{D}_{*,T^{(2)}},$ respectively. Indeed, for $h\in\mathcal{H}$ we have
$U\rho^{{}^{\prime}}(a)(\Delta_{*}(T^{(1)})h,(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}h)$
$=U(\sigma(a)\Delta_{*}(T^{(1)})h,(\phi_{1}(a)\otimes I_{\mathcal{D}_{*,T^{(2)}}})(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}h)$
$=U(\Delta_{*}(T^{(1)})\sigma(a)h,(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))(\phi_{1}(a)\otimes I_{\mathcal{H}})\widetilde{T}^{(1)^{*}}h)$
$=U(\Delta_{*}(T^{(1)})\sigma(a)h,(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}\sigma(a)h)$
$=((I_{E_{2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}}\sigma(a)h,\Delta_{*}(T^{(2)})\sigma(a)h)$
$=((\phi_{2}(a)\otimes I_{\mathcal{D}_{*,T^{(1)}}})(I_{E_{2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}}h,\sigma(a)\Delta_{*}(T^{(2)})h)$
$=\rho^{{}^{\prime\prime}}(a)U(\Delta_{*}(T^{(1)})h,(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}h)$
Moreover, if $\dim(\mathcal{D}_{*,T^{(1)}}\oplus(E_{1}\otimes\mathcal{D}_{*,T^{(2)}}))$ = $\dim((E_{2}\otimes\mathcal{D}_{*,T^{(1)}})\oplus\mathcal{D}_{*,T^{(2)}})$ which is finite, it follows that $U$ extends to a unitary module map, denoted again by $U,$ from $\mathcal{D}_{*,T^{(1)}}\oplus(E_{1}\otimes\mathcal{D}_{*,T^{(2)}})$ to $(E_{2}\otimes\mathcal{D}_{*,T^{(1)}})\oplus\mathcal{D}_{*,T^{(2)}}.$ In particular, there exists a unitary
$U:=\begin{bmatrix}A&B\\ C&\widetilde{D}\end{bmatrix}:\mathcal{D}_{*,T^{(1)}}\oplus(E_{1}\otimes\mathcal{D}_{*,T^{(2)}})\to(E_{2}\otimes\mathcal{D}_{*,T^{(1)}})\oplus\mathcal{D}_{*,T^{(2)}},$
and it satisfies Equation (2), where $A:\mathcal{D}_{*,T^{(1)}}\to E_{2}\otimes\mathcal{D}_{*,T^{(1)}},$ $B:E_{1}\otimes\mathcal{D}_{*,T^{(2)}}\to E_{2}\otimes\mathcal{D}_{*,T^{(1)}},$ $C:\mathcal{D}_{*,T^{(1)}}\to\mathcal{D}_{*,T^{(2)}}$ and $\widetilde{D}:E_{1}\otimes\mathcal{D}_{*,T^{(2)}}\to\mathcal{D}_{*,T^{(2)}}.$ For $\eta\in\mathcal{D}_{*,T^{(1)}}\oplus(E_{1}\otimes\mathcal{D}_{*,T^{(2)}})$ we have
(2.2) $U\rho^{{}^{\prime}}(a)\eta=\rho^{{}^{\prime\prime}}(a)U\eta.$
Therefore $U$ is a module map, and hence each entry of $U$ is also a module map. That is, $A,B,C$ and $\widetilde{D}$ are module maps. In particular, $A=P_{E_{2}\otimes\mathcal{D}_{*,T^{(1)}}}U|_{\mathcal{D}_{*,T^{(1)}}}:\mathcal{D}_{*,T^{(1)}}\to E_{2}\otimes\mathcal{D}_{*,T^{(1)}},$ we have $A\sigma(a)|_{\mathcal{D}_{*,T^{(1)}}}=(\phi_{2}(a)\otimes I_{\mathcal{D}_{*,T^{(1)}}})A$ for $a\in\mathcal{M}.$ That is,
$\sigma(a)|_{\mathcal{D}_{*,T^{(1)}}}A^{*}=A^{*}(\phi_{2}(a)\otimes I_{\mathcal{D}_{*,T^{(1)}}}),$
where $P_{E_{2}\otimes\mathcal{D}_{*,T^{(1)}}}$ is the orthogonal projection of $E_{2}\otimes\mathcal{H}$ onto $E_{2}\otimes\mathcal{D}_{*,T^{(1)}}.$ Indeed, by using Equation (2.2) for $h\in\mathcal{D}_{*,T^{(1)}}$ we obtain
$A\sigma(a)h=P_{E_{2}\otimes\mathcal{D}_{*,T^{(1)}}}U\sigma(a)h$ $=P_{E_{2}\otimes\mathcal{D}_{*,T^{(1)}}}(\phi_{2}(a)\otimes I_{\mathcal{D}_{*,T^{(1)}}})Uh$
$=(\phi_{2}(a)\otimes I_{\mathcal{D}_{*,T^{(1)}}})Ah.$
From Lemma 1.3, we define a completely contractive covariant representation $(\psi_{1},X)$ of $E_{2}$ on the Hilbert space $\mathcal{D}_{*,T^{(1)}}$ by
$X(\xi_{2})h:=A^{*}(\xi_{2}\otimes h),$
where $\xi_{2}\in E_{2},h\in\mathcal{D}_{*,T^{(1)}}$ and $\psi_{1}(a):=\sigma(a)|_{\mathcal{D}_{*,T^{(1)}}},a\in\mathcal{M}.$ Similarly, $(\sigma_{1},D)$ is a completely contractive covariant representation of $E_{1}$ on the Hilbert space $\mathcal{D}_{*,T^{(2)}},$ where $\sigma_{1}(a):=\sigma(a)|_{\mathcal{D}_{*,T^{(2)}}},a\in\mathcal{M}.$
The following lemma is crucial to our discussion.
Lemma 2.1.
Assume $\mathbb{E}$ to be a product system over $\mathbb{N}^{2}_{0}$. Let $(\sigma,T^{(1)},T^{(2)})$ be a completely contractive covariant representation of $\mathbb{E}$ on $\mathcal{H}$ such that $(\sigma,T^{(1)})$ is pure and $\dim(\mathcal{D}_{*,T^{(1)}}\oplus(E_{1}\otimes\mathcal{D}_{*,T^{(2)}}))$ = $\dim((E_{2}\otimes\mathcal{D}_{*,T^{(1)}})\oplus\mathcal{D}_{*,T^{(2)}})<\infty.$ Then
$(I_{E_{2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}}=A\Delta_{*}(T^{(1)})$
$\hskip 14.45377pt+\sum_{n=0}^{\infty}B(I_{E_{1}}\otimes\widetilde{D}_{n})(I_{E_{1}^{\otimes n+1}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{n+1},$
where $U=\begin{bmatrix}A&B\\ C&\widetilde{D}\end{bmatrix}$ defined as above, $\widetilde{D}_{n}=\widetilde{D}(I_{E_{1}}\otimes\widetilde{D})\dots(I_{E_{1}^{\otimes(n-1)}}\otimes\widetilde{D})$ and the series converges with respect to the strong operator topology.
Proof.
Let $h\in\mathcal{H}$ we get
$\begin{bmatrix}A&B\\ C&\widetilde{D}\end{bmatrix}\begin{bmatrix}\Delta_{*}(T^{(1)})h\\ (I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}h\end{bmatrix}=\begin{bmatrix}(I_{E_{2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}}h\\ \Delta_{*}(T^{(2)})h\end{bmatrix}.$
It follows that
(2.3) $(I_{E_{2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}}h=A\Delta_{*}(T^{(1)})h+B(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}h$
and
$\Delta_{*}(T^{(2)})h=C\Delta_{*}(T^{(1)})h+\widetilde{D}(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}h.$
This implies that
$(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}h=(I_{E_{1}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}h+$
$\quad\quad(I_{E_{1}}\otimes\widetilde{D}(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}})\widetilde{T}^{(1)^{*}}h,$
and further we get
$(I_{E_{1}^{\otimes 2}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}_{2}h$
$=(I_{E_{1}^{\otimes 2}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{2}h+(I_{E_{1}^{\otimes 2}}\otimes\widetilde{D}(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}})\widetilde{T}^{(1)^{*}}_{2}h,$
and hence
$(I_{E_{1}^{\otimes 2}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}_{2}h$
$=(I_{E_{1}^{\otimes 2}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{2}h+(I_{E_{1}^{\otimes 2}}\otimes\widetilde{D}(I_{E_{1}}\otimes\Delta_{*}(T^{(2)})))\widetilde{T}^{(1)^{*}}_{3}h.$
From Equation  (2.3) we obtain
$\begin{split}&(I_{E_{2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}}h\\ &=A\Delta_{*}(T^{(1)})h+B(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}h\\ &=A\Delta_{*}(T^{(1)})h+B(I_{E_{1}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}h\\ &\quad\quad+B(I_{E_{1}}\otimes\widetilde{D}(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}})\widetilde{T}^{(1)^{*}}h\\ &=A\Delta_{*}(T^{(1)})h+B(I_{E_{1}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}h\\ &\quad\quad+B(I_{E_{1}}\otimes\widetilde{D})(I_{E_{1}^{\otimes 2}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}_{2}h\\ &=A\Delta_{*}(T^{(1)})h+B(I_{E_{1}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}h\\ &\quad\quad+B(I_{E_{1}}\otimes\widetilde{D})(I_{E_{1}^{\otimes 2}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{2}h\\ &\quad\quad+B(I_{E_{1}}\otimes\widetilde{D})(I_{E_{1}^{\otimes 2}}\otimes\widetilde{D}(I_{E_{1}}\otimes\Delta_{*}(T^{(2)})))\widetilde{T}^{(1)^{*}}_{3}h\\ &=A\Delta_{*}(T^{(1)})h+B(I_{E_{1}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}h\\ &\quad\quad+B(I_{E_{1}}\otimes\widetilde{D})(I_{E_{1}^{\otimes 2}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{2}h\\ &\quad\quad+B(I_{E_{1}}\otimes\widetilde{D})(I_{E_{1}^{\otimes 2}}\otimes\widetilde{D})(I_{E_{1}^{\otimes 3}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}_{3}h\\ &=A\Delta_{*}(T^{(1)})h+B(I_{E_{1}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}h\\ &\quad\quad+B(I_{E_{1}}\otimes\widetilde{D})(I_{E_{1}^{\otimes 2}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{2}h\\ &\quad\quad+B(I_{E_{1}}\otimes\widetilde{D}(I_{E_{1}}\otimes\widetilde{D}))(I_{E_{1}^{\otimes 3}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}_{3}h\\ &=A\Delta_{*}(T^{(1)})h+B(I_{E_{1}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}h\\ &\quad\quad+B(I_{E_{1}}\otimes\widetilde{D})(I_{E_{1}^{\otimes 2}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{2}h\\ &\quad\quad+B(I_{E_{1}}\otimes\widetilde{D}_{2})(I_{E_{1}^{\otimes 3}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}_{3}h.\end{split}$
Therefore iteratively for each $m\in\mathbb{N}$, we have
$(I_{E_{2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}}h=$ $A\Delta_{*}(T^{(1)})h$
$+\sum_{n=0}^{m}B(I_{E_{1}}\otimes\widetilde{D}_{n})(I_{E_{1}^{\otimes n+1}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{n+1}h$
$+B(I_{E_{1}}\otimes\widetilde{D}_{m+1})(I_{E_{1}^{\otimes m+2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{m+2}h.$
Since $(\sigma,T^{(1)})$ is pure and $\|\widetilde{D}\|\leq 1,$ it gives
$\|(I_{E_{2}}\otimes$ $\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}}h-A\Delta_{*}(T^{(1)})h-$
$\sum_{n=0}^{m}B(I_{E_{1}}\otimes\widetilde{D}_{n})(I_{E_{1}^{\otimes n+1}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{n+1}h\|$
$=\|B(I_{E_{1}}\otimes\widetilde{D}_{m+1})(I_{E_{1}^{\otimes m+2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{m+2}h\|$
$\leq\|\widetilde{T}^{(1)^{*}}_{m+2}h\|\to 0~{}\mbox{as}~{}m\to\infty.$
This complets the proof. ∎
Observation 2.2.
Under the above assumptions, the conclusion of Lemma 2.1 remains valid even when we incorporate the possibility $\dim(\mathcal{D}_{*,T^{(1)}}\oplus(E_{1}\otimes\mathcal{D}_{*,T^{(2)}}))$ or $\dim((E_{2}\otimes\mathcal{D}_{*,T^{(1)}})\oplus\mathcal{D}_{*,T^{(2)}})$ is an infinite dimensional space. The conclusion of Lemma 2.1 remains valid. If $\dim(\mathcal{D}_{*,T^{(1)}}\oplus(E_{1}\otimes\mathcal{D}_{*,T^{(2)}}))=\infty$ or $\dim((E_{2}\otimes\mathcal{D}_{*,T^{(1)}})\oplus\mathcal{D}_{*,T^{(2)}})=\infty,$ then there exists an infinite dimensional Hilbert space $\mathcal{D}$ such that the isometry
$U:\{\Delta_{*}(T^{(1)})h\oplus(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}h:~{}h\in\mathcal{H}\}\oplus\{0_{\mathcal{D}}\}\to$
$\>\>\>\>\>\>\>\>\{(I_{E_{2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}}h\oplus\Delta_{*}(T^{(2)})h:~{}h\in\mathcal{H}\}\oplus\{0_{\mathcal{D}}\}$
is defined by
$U(\Delta_{*}(T^{(1)})h,(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}h,0_{\mathcal{D}})$
$=((I_{E_{2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}}h,\Delta_{*}(T^{(2)})h,0_{\mathcal{D}}),$
for all $h\in\mathcal{H}$ extends to a unitary, denoted again by $U,$ from $\mathcal{D}_{*,T^{(1)}}\oplus(E_{1}\otimes\mathcal{D}_{*,T^{(2)}})\oplus\mathcal{D}$ to $(E_{2}\otimes\mathcal{D}_{*,T^{(1)}})\oplus\mathcal{D}_{*,T^{(2)}}\oplus\mathcal{D}.$ Now we proceed similarly with the unitary matrix
$U=\begin{bmatrix}A&B\\ C&\widetilde{D}\end{bmatrix}:\mathcal{D}_{*,T^{(1)}}\oplus(E_{1}\otimes\mathcal{D}_{*,T^{(2)}})\oplus\mathcal{D}\to(E_{2}\otimes\mathcal{D}_{*,T^{(1)}})\oplus\mathcal{D}_{*,T^{(2)}}\oplus\mathcal{D},$
defined by
$U(\Delta_{*}(T^{(1)})h,(I_{E_{1}}\otimes\Delta_{*}(T^{(2)}))\widetilde{T}^{(1)^{*}}h,0_{\mathcal{D}})$
$=((I_{E_{2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}}h,\Delta_{*}(T^{(2)})h,0_{\mathcal{D}}),$
to obtain the same conclusion as in Lemma 2.1.
2.1. Transfer Function
Suppose that $\mathbb{E}$ is a product system over $\mathbb{N}_{0}^{2},$ and $\pi$ is a representation of $\mathcal{M}$ on $\mathcal{K}.$ Let $\Theta:E_{2}\longrightarrow B(\mathcal{K},\mathcal{F}(E_{1})\otimes_{\pi}\mathcal{K})$ be a completely bounded bi-module map. Define a bounded linear map $\widetilde{\Theta}:E_{2}\otimes\mathcal{K}\longrightarrow\mathcal{F}(E_{1})\otimes_{\pi}\mathcal{K}$ by $\widetilde{\Theta}(\xi\otimes h)=\Theta(\xi)h$ for all $\xi\in E_{2},h\in\mathcal{K},$ and it satisfies $\widetilde{\Theta}(\phi_{2}(a)\otimes I_{\mathcal{K}})=\rho(a)\widetilde{\Theta},a\in\mathcal{M},$ where $\phi_{2}$ is the left action of $\mathcal{M}$ on $E_{2}.$ We define a corresponding completely bounded bi-module map [20] $M_{\Theta}:E_{2}\longrightarrow B(\mathcal{F}(E_{1})\otimes_{\pi}\mathcal{K})$ by
$M_{\Theta}(\xi)(S_{n}^{(1)}(\xi_{n})h)=\widetilde{S}_{n}^{(1)}(I_{E_{1}^{\otimes n}}\otimes\widetilde{\Theta})(t_{2,1}^{(1,n)}\otimes I_{\mathcal{K}})(\xi\otimes\xi_{n}\otimes h),$
[…]
$M_{\Theta}(\xi)\left(\bigoplus_{n\in\mathbb{N}_{0}}\xi_{n}\otimes h_{n}\right)=\sum_{n\in\mathbb{N}_{0}}\widetilde{S}_{n}^{(1)}(I_{E_{1}^{\otimes n}}\otimes\widetilde{\Theta})(t_{2,1}^{(1,n)}\otimes I_{\mathcal{K}})(\xi\otimes\xi_{n}\otimes h_{n}),$
where $\xi\in E_{2},\xi_{n}\in E_{1}^{\otimes n},h_{n}\in\mathcal{K}.$ Let $(\rho,S)$ be the induced representation of $E_{1}$ induced by $\pi,$ then $(\rho,S,M_{\Theta})$ is a c.b.c. representation of $\mathbb{E}$ on $\mathcal{F}(E_{1})\otimes_{\pi}\mathcal{K}.$ Note that $(\rho,M_{\Theta})$ is an isometric covariant representation if and only if $\widetilde{\Theta}$ is an isometry.
With the help of unitary module map $U$ as we defined above, we define a bounded linear map $\widetilde{\Theta}:E_{2}\otimes\mathcal{D}_{*,T^{(1)}}\to F(E_{1})\otimes\mathcal{D}_{*,T^{(1)}}$ by
$\widetilde{\Theta}:=A^{*}\oplus\bigoplus_{n=0}^{\infty}(I_{E_{1}^{\otimes n+1}}\otimes C^{*})(I_{E_{1}}\otimes\widetilde{D}^{*}_{n})B^{*}.$
First we will show that it is well defined, it is enough to prove that $\|\bigoplus_{n=0}^{\infty}(I_{E_{1}^{\otimes n+1}}\otimes C^{*})(I_{E_{1}}\otimes\widetilde{D}^{*}_{n})B^{*}(\eta)\|^{2}<\infty$ for $\eta\in E_{2}\otimes\mathcal{D}_{*,T^{(1)}}.$ Indeed, using $UU^{*}=I$ we have
$\|\bigoplus_{n=0}^{\infty}(I_{E_{1}^{\otimes n+1}}\otimes C^{*})(I_{E_{1}}\otimes\widetilde{D}^{*}_{n})B^{*}(\eta)\|^{2}$
$=\sum_{n=0}^{\infty}\langle B(I_{E_{1}}\otimes\widetilde{D}_{n})(I_{E_{1}^{\otimes n+1}}\otimes CC^{*})(I_{E_{1}}\otimes\widetilde{D}^{*}_{n})B^{*}\eta,\eta\rangle$
$=\sum_{n=0}^{\infty}\langle B(I_{E_{1}}\otimes\widetilde{D}_{n})(I_{E_{1}^{\otimes n+1}}\otimes(I_{\mathcal{D}_{*,T^{(2)}}}-\widetilde{D}\widetilde{D}^{*}))(I_{E_{1}}\otimes\widetilde{D}^{*}_{n})B^{*}\eta,\eta\rangle$
$=\sum_{n=0}^{\infty}\langle\big{(}B(I_{E_{1}}\otimes\widetilde{D}_{n}\widetilde{D}^{*}_{n})B^{*}-B(I_{E_{1}}\otimes\widetilde{D}_{n+1}\widetilde{D}_{n+1}^{*})B^{*}\big{)}\eta,\eta\rangle$
$=\langle BB^{*}-\lim_{n\rightarrow\infty}B(I_{E_{1}}\otimes\widetilde{D}_{n}\widetilde{D}^{*}_{n})B^{*}\eta,\eta\rangle$
$=\|B^{*}\eta\|^{2}-\lim_{n\rightarrow\infty}\langle B(I_{E_{1}}\otimes\widetilde{D}_{n}\widetilde{D}^{*}_{n})B^{*}\eta,\eta\rangle<\|B^{*}\eta\|^{2}.$
Here the last inequality follows by the fact that $B$ and $\widetilde{D}$ are contractions. This shows that $\widetilde{\Theta}$ is well defined and contractive.
Since $(\sigma_{1},D)$ is a completely contractive covariant representation of $E_{1}$ on $\mathcal{D}_{*,T^{(2)}},$ from Theorem 1.5, $\mathcal{D}_{*,T^{(2)}}=\mathcal{D}^{(1)}_{*,T^{(2)}}\oplus\mathcal{D}^{(2)}_{*,T^{(2)}}$ has unique orthogonal decomposition such that $\mathcal{D}^{(1)}_{*,T^{(2)}},$ $\mathcal{D}^{(2)}_{*,T^{(2)}}$ reduces $(\sigma_{1},D),$ $(\pi_{1},V^{(1)}):=(\sigma_{1},D)|_{\mathcal{D}^{(1)}_{*,T^{(2)}}}$ is an isometric as well as co-isometric covariant representation and $(\pi_{2},V^{(2)}):=(\sigma_{1},D)|_{\mathcal{D}^{(2)}_{*,T^{(2)}}}$ is a completely non-unitary covariant representation. That is, $\widetilde{D}^{*}\widetilde{D}|_{E_{1}\otimes\mathcal{D}^{(1)}_{*,T^{(2)}}}=I_{E_{1}\otimes\mathcal{D}^{(1)}_{*,T^{(2)}}}.$ Also using the fact that $U$ is unitary, we have $B^{*}B+\widetilde{D}^{*}\widetilde{D}=I_{E_{1}\otimes\mathcal{D}_{*,T^{(2)}}},$ therefore $B^{*}B|_{E_{1}\otimes\mathcal{D}^{(1)}_{*,T^{(2)}}}=0,$ i.e., $E_{1}\otimes\mathcal{D}^{(1)}_{*,T^{(2)}}\subseteq ker(B),$ or equivalently, ${Im(B^{*})}\subseteq E_{1}\otimes\mathcal{D}^{(2)}_{*,T^{(2)}}.$ Hence $(I_{E_{1}}\otimes\widetilde{D}_{n}\widetilde{D}_{n}^{*})B^{*}=(I_{E_{1}}\otimes\widetilde{V}_{n}^{(2)}\widetilde{V}_{n}^{(2)*})B^{*}.$ Finally, we can rewrite $\lim_{n\rightarrow\infty}B(I_{E_{1}}\otimes\widetilde{D}_{n}\widetilde{D}_{n}^{*})B^{*}=\lim_{n\rightarrow\infty}B(I_{E_{1}}\otimes\widetilde{V}_{n}^{(2)}\widetilde{V}_{n}^{(2)*})B^{*}.$
Observation 2.3.
$\widetilde{\Theta}$ is an isometry if and only if $(\pi_{2},V^{(2)})$ is pure.
Define ${\Theta}:E_{2}\to B(\mathcal{D}_{*,T^{(1)}},F(E_{1})\otimes\mathcal{D}_{*,T^{(1)}})$ by ${\Theta}(\eta)h=\widetilde{\Theta}(\eta\otimes h),$ where $\eta\in E_{2}$ and $h\in\mathcal{D}_{*,T^{(1)}}.$ We need to show that $\Theta$ is a bi-module map. That is, $\Theta(a.\eta.b)=\rho(a)\Theta(\eta)\rho(b).$ Let $h\in\mathcal{D}_{*,T^{(1)}}$ we have
$\widetilde{\Theta}(a.\eta.b\otimes h)=A^{*}(a.\eta.b\otimes h)\oplus\bigoplus_{n=0}^{\infty}(I_{E_{1}^{\otimes n+1}}\otimes C^{*})(I_{E_{1}}\otimes\widetilde{D}^{*}_{n})B^{*}(a.\eta.b\otimes h)$
$=A^{*}(a.\eta\otimes b.h)\oplus\bigoplus_{n=0}^{\infty}(I_{E_{1}^{\otimes n+1}}\otimes C^{*})(I_{E_{1}}\otimes\widetilde{D}^{*}_{n})B^{*}(a.\eta\otimes b.h)$
$=\big{(}A^{*}(\phi_{2}(a)\otimes I_{\mathcal{D}_{*,T^{(1)}}})\oplus$
$\quad\quad\bigoplus_{n=0}^{\infty}(I_{E_{1}}\otimes(I_{E_{1}^{\otimes n}}\otimes C^{*})\widetilde{D}^{*}_{n})B^{*}(\phi_{2}(a)\otimes I_{\mathcal{D}_{*,T^{(1)}}})\big{)}(\eta\otimes b.h)$
$=\sigma(a)|_{\mathcal{D}_{*,T^{(1)}}}A^{*}(\eta\otimes b.h)\oplus$
$\quad\quad\bigoplus_{n=0}^{\infty}(I_{E_{1}^{\otimes n+1}}\otimes C^{*})(I_{E_{1}}\otimes\widetilde{D}^{*}_{n})(\phi_{1}(a)\otimes I_{\mathcal{D}_{*,T^{(2)}}})B^{*}(\eta\otimes b.h)$
$=\sigma(a)|_{\mathcal{D}_{*,T^{(1)}}}A^{*}(\eta\otimes b.h)\oplus$
$\quad\quad\bigoplus_{n=0}^{\infty}(I_{E_{1}^{\otimes n+1}}\otimes C^{*})(\phi_{1}^{n+1}(a)\otimes I_{\mathcal{D}_{*,T^{(2)}}})(I_{E_{1}}\otimes\widetilde{D}_{n}^{*})B^{*}(\eta\otimes b.h)$
$=\sigma(a)|_{\mathcal{D}_{*,T^{(1)}}}A^{*}(\eta\otimes b.h)\oplus$
$\quad\quad\bigoplus_{n=0}^{\infty}(\phi_{1}^{n+1}(a)\otimes I_{\mathcal{D}_{*,T^{(1)}}})(I_{E_{1}^{\otimes n+1}}\otimes C^{*})(I_{E_{1}}\otimes\widetilde{D}_{n}^{*})B^{*}(\eta\otimes b.h).$
Finally, we obtain
$\widetilde{\Theta}(a.\eta.b\otimes h)=\big{(}\sigma(a)|_{\mathcal{D}_{*,T^{(1)}}}A^{*}\oplus$
$\quad\quad\bigoplus_{n=0}^{\infty}(\phi_{1}^{n+1}(a)\otimes I_{\mathcal{D}_{*,T^{(1)}}})(I_{E_{1}}\otimes(I_{E_{1}^{\otimes n}}\otimes C^{*})\widetilde{D}_{n}^{*})B^{*}\big{)}(\eta.b\otimes h)$
$=\big{(}\big{(}\sigma(a)|_{\mathcal{D}_{*,T^{(1)}}}\oplus\bigoplus_{n=0}^{\infty}(\phi_{1}^{n+1}(a)\otimes I_{\mathcal{D}_{*,T^{(1)}}})\big{)}(A^{*}\oplus$
$\quad\quad\bigoplus_{n=0}^{\infty}(I_{E_{1}}\otimes(I_{E_{1}^{\otimes n}}\otimes C^{*})\widetilde{D}_{n}^{*})B^{*}))\big{(}\eta.b\otimes h)$
$=\rho(a)\Theta(\eta)\rho(b)h.$
Thus $\Theta$ is a bi-module map. From [20, Lemma 2.2], $(\rho,S^{(1)},M_{\Theta})$ is a completely contractive covariant representation of $\mathbb{E}$ over $\mathbb{N}_{0}^{2}$ on $\mathcal{F}(E_{1})\otimes\mathcal{D}_{*,T^{(1)}}.$ The contractive linear map $\widetilde{\Theta}$ is called the transfer function of unitary module map $U.$
2.2. Ando Dilations
We are ready to prove our main result, a generalization of [5, Theorem 3.1].
Theorem 2.4.
Assume $\mathbb{E}$ to be a product system over $\mathbb{N}^{2}_{0}$. Let $(\sigma,T^{(1)},T^{(2)})$ be a completely contractive covariant representation of $\mathbb{E}$ on $\mathcal{H}$ such that $(\sigma,T^{(1)})$ is pure and $\dim(\mathcal{D}_{*,T^{(1)}}\oplus(E_{1}\otimes\mathcal{D}_{*,T^{(2)}}))$ = $\dim((E_{2}\otimes\mathcal{D}_{*,T^{(1)}})\oplus\mathcal{D}_{*,T^{(2)}})<\infty.$ Then there exist an isometry $\Pi_{1}:\mathcal{H}\longrightarrow\mathcal{F}(E_{1})\otimes\mathcal{D}_{*,T^{(1)}}$ and a completely contractive bi-module map ${\Theta}:E_{2}\to B(\mathcal{D}_{*,T^{(1)}},F(E_{1})\otimes\mathcal{D}_{*,T^{(1)}})$ such that
$\Pi_{1}\sigma(a)=\rho(a)\Pi_{1},(I_{E_{1}}\otimes\Pi_{1})\widetilde{T}^{(1)*}=\widetilde{S}^{(1)*}\Pi_{1}\>\>\mbox{and}\>\>{(I_{E_{2}}\otimes\Pi_{1})\widetilde{T}^{(2)}}^{*}=\widetilde{M}^{*}_{\Theta}\Pi_{1},$
for all $a\in\mathcal{M}.$
Proof.
Let $(\sigma,T^{(1)})$ be a pure completely contractive covariant representation of $E_{1}$ on $\mathcal{H},$ then from Theorem 1.6, there exists an isometry $\Pi_{1}:\mathcal{H}\longrightarrow\mathcal{F}(E_{1})\otimes\mathcal{D}_{*,T^{(1)}}$ defined by
$\Pi_{1}h=\sum_{k\geq 0}(I_{E_{1}^{\otimes k}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{k}h,\>\>\>h\in\mathcal{H}$
such that
$\Pi_{1}\sigma(a)=\rho(a)\Pi_{1}\>\>\mbox{and}\>\>(I_{E_{1}}\otimes\Pi_{1})\widetilde{T}^{(1)*}=\widetilde{S}^{(1)*}\Pi_{1},$
where $(\rho,S^{(1)})$ is an induced representation of $E_{1}$ induced by $\sigma|_{\mathcal{D}_{*,T^{(1)}}}.$ We only need to show that ${(I_{E_{2}}\otimes\Pi_{1})\widetilde{T}^{(2)}}^{*}=\widetilde{M}^{*}_{\Theta}\Pi_{1}.$ Define a completely contractive bi-module map ${\Theta}:E_{2}\to B(\mathcal{D}_{*,T^{(1)}},F(E_{1})\otimes\mathcal{D}_{*,T^{(1)}})$ by $\Theta(\xi)h=\widetilde{\Theta}(\xi\otimes h)$ where $\xi\in E_{2},h\in\mathcal{D}_{*,T^{(1)}}$ and $\widetilde{\Theta}=A^{*}\oplus\bigoplus_{m=0}^{\infty}(I_{E_{1}^{\otimes m+1}}\otimes C^{*})(I_{E_{1}}\otimes\widetilde{D}_{m}^{*})B^{*}.$ Define a corresponding contractive linear map $\widetilde{M}_{\Theta}:E_{2}\otimes F(E_{1})\otimes\mathcal{D}_{*,T^{(1)}}\longrightarrow F(E_{1})\otimes\mathcal{D}_{*,T^{(1)}}$ by
$\widetilde{M}_{\Theta}=\widetilde{S}^{(1)}_{n}(I_{E_{1}^{\otimes n}}\otimes\widetilde{\Theta})(t_{2,1}^{(1,n)}\otimes I_{\mathcal{D}_{*,T^{(1)}}}),$
and it satisfies $\widetilde{M}_{\Theta}(\phi_{2}(a)\otimes I_{F(E_{1})\otimes\mathcal{D}_{*,T^{(1)}}})=\rho(a)\widetilde{M}_{\Theta}$ for $a\in\mathcal{M}.$ For $\eta_{n}\in{E_{2}}\otimes E_{1}^{\otimes n}\otimes\mathcal{D}_{*,T^{(1)}},h\in\mathcal{H}$ and from Lemma 2.1 we have
$\langle\widetilde{M}^{*}_{\Theta}\Pi_{1}h,\eta_{n}\rangle$
$=\langle\sum_{k=0}^{\infty}(I_{E_{1}^{\otimes k}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{k}h,\widetilde{S}_{n}^{(1)}(I_{E_{1}^{\otimes n}}\otimes\widetilde{\Theta})(t_{2,1}^{(1,n)}\otimes I_{\mathcal{D}_{*,T^{(1)}}})\eta_{n}\rangle$
$=\langle(I_{E_{1}^{\otimes n}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{n}h,(I_{E_{1}^{\otimes n}}\otimes A^{*})(t_{2,1}^{(1,n)}\otimes I_{\mathcal{D}_{*,T^{(1)}}})\eta_{n}\rangle$
$\quad\quad+\sum_{m=0}^{\infty}\langle(I_{E_{1}^{\otimes n+m+1}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{n+m+1}h,$
$\quad\quad\quad(I_{E_{1}^{\otimes n}}\otimes(I_{E_{1}}\otimes(I_{E_{1}^{\otimes m}}\otimes C^{*})\widetilde{D}_{m}^{*})B^{*})(t_{2,1}^{(1,n)}\otimes I_{\mathcal{D}_{*,T^{(1)}}})\eta_{n}\rangle$
$=\langle(I_{E_{1}^{\otimes n}}\otimes A\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{n}h,(t_{2,1}^{(1,n)}\otimes I_{\mathcal{D}_{*,T^{(1)}}})\eta_{n}\rangle$
$\quad\quad+\sum_{m=0}^{\infty}\langle(I_{E_{1}^{\otimes n}}\otimes(B(I_{E_{1}}\otimes\widetilde{D}_{m})))(I_{E_{1}^{\otimes n+m+1}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{n+m+1}h,$
$\quad\quad\quad(t_{2,1}^{(1,n)}\otimes I_{\mathcal{D}_{*,T^{(1)}}})\eta_{n}\rangle$
$=\langle\big{(}I_{E_{1}^{\otimes n}}\otimes A\Delta_{*}(T^{(1)})+\sum_{m=0}^{\infty}(I_{E_{1}^{\otimes n}}\otimes(B(I_{E_{1}}\otimes\widetilde{D}_{m})))$
$\quad\quad(I_{E_{1}^{\otimes n+m+1}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{m+1})\big{)}\widetilde{T}^{(1)^{*}}_{n}h,(t_{2,1}^{(1,n)}\otimes I_{\mathcal{D}_{*,T^{(1)}}})\eta_{n}\rangle$
$=\langle(I_{E_{1}^{\otimes n}}\otimes\big{(}A\Delta_{*}(T^{(1)})+\sum_{m=0}^{\infty}B(I_{E_{1}}\otimes\widetilde{D}_{m})(I_{E_{1}^{\otimes m+1}}\otimes C\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{m+1})\big{)}\widetilde{T}^{(1)^{*}}_{n}h,$
$\quad\quad(t_{2,1}^{(1,n)}\otimes I_{\mathcal{D}_{*,T^{(1)}}})\eta_{n}\rangle$
$=\langle(I_{E_{1}^{\otimes n}}\otimes(I_{E_{2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}}\big{)}\widetilde{T}^{(1)^{*}}_{n}h,(t_{2,1}^{(1,n)}\otimes I_{\mathcal{D}_{*,T^{(1)}}})\eta_{n}\rangle.$
On the other hand, since
$\langle{(I_{E_{2}}\otimes\Pi_{1})\widetilde{T}^{(2)}}^{*}h,\eta_{n}\rangle=\langle{(I_{E_{2}}\otimes\sum_{k\geq 0}(I_{E_{1}^{\otimes k}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{k})\widetilde{T}^{(2)}}^{*}h,\eta_{n}\rangle$
$=\langle{(I_{E_{2}}\otimes(I_{E_{1}^{\otimes n}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(1)^{*}}_{n})\widetilde{T}^{(2)}}^{*}h,\eta_{n}\rangle$
$=\langle h,{\widetilde{T}^{(2)}(I_{E_{2}}\otimes\widetilde{T}^{(1)}_{n}(I_{E_{1}^{\otimes n}}\otimes\Delta_{*}(T^{(1)})))}\eta_{n}\rangle$
$=\langle h,{\widetilde{T}^{(1)}_{n}(I_{E_{1}^{\otimes n}}\otimes\widetilde{T}^{(2)}(I_{E_{2}}\otimes\Delta_{*}(T^{(1)})))}(t_{2,1}^{(1,n)}\otimes I_{\mathcal{D}_{*,T^{(1)}}})\eta_{n}\rangle$
$=\langle(I_{E_{1}^{\otimes n}}\otimes(I_{E_{2}}\otimes\Delta_{*}(T^{(1)}))\widetilde{T}^{(2)^{*}})\widetilde{T}^{(1)^{*}}_{n}h,(t_{2,1}^{(1,n)}\otimes I_{\mathcal{D}_{*,T^{(1)}}})\eta_{n}\rangle=\langle\widetilde{M}^{*}_{\Theta}\Pi_{1}h,\eta_{n}\rangle,$
we get ${(I_{E_{2}}\otimes\Pi_{1})\widetilde{T}^{(2)}}^{*}=\widetilde{M}^{*}_{\Theta}\Pi_{1}.$ This complets the proof. ∎
Observation 2.5.
  1. (1)
    In the setting of Theorem 2.4, if $\mathcal{Q}:=\Pi_{1}\mathcal{H},$ then
    $\rho(a)\mathcal{Q}\subseteq\mathcal{Q},~{}\widetilde{S}^{(1)^{*}}(\mathcal{Q})\subseteq E_{1}\otimes\mathcal{Q}\quad\mbox{and}\quad\widetilde{M}_{\Theta}^{*}(\mathcal{Q})\subseteq E_{2}\otimes\mathcal{Q}\quad\mbox{for}\quad a\in\mathcal{M}.$
  2. (2)
    The representations $(\pi^{\prime},X,Y)$ and $(\sigma,T^{(1)},T^{(2)})$ of $\mathbb{E}$ acting on the Hilbert spaces $\mathcal{Q}$ and $\mathcal{H},$ respectively, are isomorphic, where $\pi^{\prime}=\rho|_{\mathcal{Q}},$ $\widetilde{X}=P_{\mathcal{Q}}\widetilde{S}^{(1)}|_{E_{1}\otimes\mathcal{Q}}$ and $\widetilde{Y}=P_{\mathcal{Q}}\widetilde{M}_{\Theta}|_{E_{2}\otimes\mathcal{Q}}.$
  3. (3)
    Theorem 2.4 remains valid if we drop the assumption that $\dim(\mathcal{D}_{*,T^{(1)}}\oplus(E_{1}\otimes\mathcal{D}_{*,T^{(2)}}))<\infty$ and $\dim((E_{2}\otimes\mathcal{D}_{*,T^{(1)}})\oplus\mathcal{D}_{*,T^{(2)}})<\infty.$
Conflict of Interest
We state that there is no conflict of interest and we have no personal relationships that could have appeared to influence the work reported in this paper.
Data Availability
Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.
Acknowledgement
The authors thank the reviewer for carefully reading the manuscript and the editor for suggesting changes in the manuscript. The authors want to thank Harsh Trivedi and Shankar Veerabathiran for some fruitful discussions.